OpenStudy (anonymous):

During a paint sale, a hardware store sold 33 gallons of flat paint and 57 gallons of high-gloss paint. What is the ratio of gallons of high-gloss sold to the total gallons sold????....please help

OpenStudy (anonymous):

57/90

OpenStudy (anonymous):

57/(33+57) = 57/90 =19/30